OpenStudy (anonymous):

Solve: –(3x – 2) = –4(x + 1) – 4

OpenStudy (anonymous):

-3x+2=-4x-4-4 -3x+2=-4x-8 -10=x

OpenStudy (anonymous):

-3x+2=-4x-8